A car starts from rest after 5 seconds and it accelerates uniformly for 10 seconds before attaining a speed of 20m/s .If it maintains this speed for the next 5 seconds draw a sketch to describe the journey, use the sketch to determine the uniform acceleration, the total distance covered,the average velocity.
1
Expert's answer
2019-11-11T15:52:58-0500


The uniform acceleration:


"a=\frac{20-0}{15-5}=2\frac{m}{s^2}"

The


"v=\frac{0.5(20)(15+5)}{20}=10\frac{m}{s}"

The total distance covered:


"s=0.5(20)(15+5)=200\ m"